# Neutrino masses, matter-antimatter asymmetry, dark matter, and supermassive black hole formation explained with Majorons

### v1

@@ I Introduction
There are a number of important questions which are unadressed by the otherwise-beautiful Standard Model of particle physics (SM). These include the origin of neutrino mass, the production of the observed baryon asymmetry of the Universe, and the nature of dark matter. One well-motivated solution to the origin of the neutrino masses involves extending the Standard Model to include a spontaneously broken global lepton number symmetry $U(1)_{L}$. If the lepton number symmetry is broken by the vacuum expectation value (VEV) of a singlet scalar at a high scale, right-handed neutrinos acquire large Majorana masses, which can ‘seesaw’ the left-handed neutrinos, forcing them to acquire naturally light masses [1, 2, 3, 4, 5]. In the theory of leptogenesis [6, 7], the CP-violating decay of these right handed-neutrinos in the early Universe is then able to generate a lepton asymmetry. This is then converted to a baryon asymmetry through sphaleron processes [8]. The angular mode of the singlet scalar gives rise to a pseudo Nambu-Goldstone boson known as the Majoron [9, 10, 11].
In this unified framework, it is even more appealing to consider the possibility that the Majoron is a dark matter candidate [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. When the Majoron is heavy ($>\mathrm{keV}$) and interacts weakly or feebly, it can be produced via freeze-out or freeze-in [16]. Such a scenario could be probed by the emission of neutrino lines, X-rays and gamma-rays [20, 21, 22, 23, 24, 25, 26].
@@
$U(1)_{EM}$ $(2/3,-1/3)$ $2/3$ $-1/3$ $(0,-1)$ $-1$ 00 $(1,0)$ $(1,0)$ 00
$U(1)_{L}$ $\mathcal{X}_{q}$ $\mathcal{X}_{q}+\mathcal{X}_{1}$ $\mathcal{X}_{q}-\mathcal{X}_{1}$ $1/2-\mathcal{X}_{1}$ $1/2-\mathcal{X}_{1}-\mathcal{X}_{2}$ $1/2$ $\mathcal{X}_{1}$ $\mathcal{X}_{2}$ $1$
@@ II The Two Higgs Doublet Majoron Model
It is worth noting that the quark sector only couples to $H_{1}$, and the charged leptons and the neutrinos separately couple to $H_{2}$ and $H_{1}$. Thus, this model does not induce flavor changing neutral currents (FCNCs). The Yukawa couplings do not entirely fix the lepton numbers. Once the U$(1)_{L}$ charge of $\phi$ is chosen to be $\mathcal{X}_{\phi}=1$, the remaining degrees of freedom in the assignment of $U(1)_{L}$ charge can be parametrized in the charges of left-handed quarks ($\mathcal{X}_{q}$), and the charges of the two Higgs ($\mathcal{X}_{1}$ and $\mathcal{X}_{2}$). The electromagnetic and the general $U(1)_{L}$ charge assignment of the model are outlined in Table 1.
@@
$=n_{f}\left(\mathcal{X}_{2}-\mathcal{X}_{1}\right)$
Similarly, one can also show that this model does not have a QCD anomaly, regardless again of the choices for $\mathcal{X}_{q},\mathcal{X}_{1}$ and $\mathcal{X}_{2}$. Here, we consider a general charge assignment for $\mathcal{X}_{2}-\mathcal{X}_{1}\equiv\mathcal{X}_{12}$. Cases of particular interest include $\mathcal{X}_{12}=1~{}\mathrm{or}~{}2$, corresponding to the model in Ref. [27] and the more generally case $\mathcal{X}_{12}>2$. In this latter case, the Higgs doublets are charged from a high dimensional operator of the form
@@
[…]
@@
The additional couplings $|\phi|^{2}\left|H_{1}\right|^{2}$ and $|\phi|^{2}\left|H_{2}\right|^{2}$ in Eq. (7) may spoil the hierarchy $v_{\phi}\gg v_{EW}$ if their coefficients are of $\mathcal{O}(1)$. However, it is argued in Refs. [64] and [65] that it is natural to set $\lambda_{\phi 1}$ and $\lambda_{\phi 2}$ to be small, since if they vanish, we obtain an enhanced Poincare symmetry for the model. Therefore, the Higgs sector remains radiatively stable even with the spontaneously broken $U(1)_{L}$ at a much higher scale  [^1] Without $V_{\mathrm{mix}}$, the scalar sector has a larger symmetry which corresponds to independent rotations of $H_{1}$, $H_{2}$ and $\phi$. If the theory is close to a conformal fixed point below the lepton number breaking scale, the two Higgs doublets are almost massless under this larger symmetry group. The hierarchy $v_{\phi}\gg v_{EW}$ can be equivalently maintained in this way, and the electroweak scale is set by $V_{\mathrm{mix}}$ where $\mu_{12}\sim\mathrm{TeV}$ can be achieved with suitable choices of $\Lambda$ and $v_{\phi}$..
@@ III Majoron Dark Matter at eV Scales
We note that when Majoron comprise the dark matter, the previously-discussed natural choice of $F_{J}\gtrsim 10^{11}$eV corresponds nicely to Majorons in the eV mass range. Beyond the aforementioned theoretical motivations, the eV mass range is important also for practical reasons: 1) using existing telescope observations, we can probe a significant fraction of the parameter space for the Majoron-to-photon decay signals and constrain our Majoron models with $\mathcal{X}_{12}>2$, and 2) for Majorons that can decay into Lyman-Werner photons, the supplied radiation background can be strong enough to trigger the formation of direct collapse black holes [32]. Therefore, this eV mass window could connect the problem of early SMBH formation to the nature of neutrino mass and dark matter, while laying within reach of current instruments.
@@ IV SMBH from Majoron Decay
One of the most striking discovery from the JWST is the surprising abundance of active galactic nuclei (AGNs) powered by SMBHs at high redshifts [43, 44, 45, 46, 47, 48, 49, 50, 51]. The existence of this population of seemingly overmassive black holes challenges our understanding of their origin and growth history. Producing such early SMBHs from ‘light seeds’ (e.g. the remnants of Pop III stars) requires extended periods of super-Eddington accretion [82, 83, 84], but even assuming this is possible, feedback events could stall subsequent accretion [85, 86] and it remains challenging to rapidly grow SMBHs from this route. On the other hand, direct collapse black holes [34, 35, 36, 37, 38, 39, 40, 41, 42] can provide heavy seeds for SMBH growth, being born as black holes at masses closer to $10^{5}M_{\odot}$. Although either seed scenarios cannot be excluded by the current JWST data [87, 88], SMBHs discovered in some extreme environments [89, 90, 91] may favor the direct collapse scenario.
@@
$k_{\mathrm{H}_{2}}(z)\approx 1.39\times 10^{-12}\mathrm{~{}s}^{-1}\frac{J_{LW}(z)}{J_{21}}.$
@@
Due to the in-situ production of radiation, the self-shielding effect should be greatly suppressed and we expect $\varepsilon_{sh}\ll 1$. We use the same chemical network as in Ref. [32] to track the change in temperature and chemical components during the collapse. A successful direct collapse process is identified by a final temperature around $10^{4}~{}\text{K}$ and an $\mathrm{H_{2}}$ fraction $\ll 10^{-5}$.

### v2

@@ I Introduction
A number of important questions are not explained by the Standard Model of particle physics (SM). These include the origin of neutrino mass, the production of the observed baryon asymmetry of the Universe, and the nature of dark matter. One well-motivated solution to the origin of the neutrino masses involves extending the Standard Model to include a spontaneously broken global lepton number symmetry $U(1)_{L}$. If the lepton number symmetry is broken by the vacuum expectation value (VEV) of a singlet scalar at a high scale, right-handed neutrinos acquire large Majorana masses, which can ‘seesaw’ the left-handed neutrinos, forcing them to acquire naturally light masses [1, 2, 3, 4, 5]. In the theory of leptogenesis [6, 7], the CP-violating decay of these right handed-neutrinos in the early Universe is then able to generate a lepton asymmetry. This is then converted to a baryon asymmetry through sphaleron processes [8]. The angular mode of the singlet scalar gives rise to a pseudo Nambu-Goldstone boson known as the Majoron [9, 10, 11].
In this unified framework, it is even more appealing to consider the possibility that the Majoron is a dark matter candidate [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. When the Majoron is heavy ($\mathrm{keV}$) and interacts weakly or feebly, it can be produced via freeze-out or freeze-in [16]. Such a scenario could be probed by the emission of neutrino lines, X-rays and gamma-rays [20, 21, 22, 23, 24, 26, 27].
@@
$U(1)_{EM}$ (2/3,1/3)(2/3,-1/3) $2/3$ 1/3-1/3 (0,1)(0,-1) 1-1 0 $(1,0)$ $(1,0)$ 0
$U(1)_{L}$ $\mathcal{X}_{q}$ $\mathcal{X}_{q}+\mathcal{X}_{1}$ $\mathcal{X}_{q}-\mathcal{X}_{1}$ $1/2-\mathcal{X}_{1}$ $\mathcal{X}_{1}-\mathcal{X}_{2} 1/21/2$ $\mathcal{X}_{1}$ $\mathcal{X}_{2}$ 11
@@ II The Two Higgs Doublet Majoron Model
[…]
@@
$=n_{f}\left(\mathcal{X}_{2}-\mathcal{X}_{1}\right)~.$
Similarly, one can also show that this model does not have a QCD anomaly, regardless again of the choices for $\mathcal{X}_{q},\mathcal{X}_{1}$ and $\mathcal{X}_{2}$. Here, we consider a general charge assignment for $\mathcal{X}_{2}-\mathcal{X}_{1}\equiv\mathcal{X}_{12}$. Cases of particular interest include $\mathcal{X}_{12}=1~\mathrm{or}~2$, corresponding to the model in Ref. [28] and the more general case $\mathcal{X}_{12}>2$. As we will see, we will generally be interested in larger electromagnetic anomalies in order to successfully explain direct collapse SMBH formation, and so we will focus on the case $\mathcal{X}_{12}>2$. In this latter case, the Higgs doublets are charged from a high dimensional operator of the form
@@
$V_{\mathrm{mix}}\supset\frac{\lambda_{\phi 12}v_{\phi}^{n}}{2^{n/2}\Lambda^{n-2}}H_{2}^{\dagger}H_{1}+\text{ h.c. }\equiv\mu_{12}^{2}H_{2}^{\dagger}H_{1}+\text{ h.c. }.$ (8)
@@
The additional couplings $|\phi|^{2}\left|H_{1}\right|^{2}$ and $|\phi|^{2}\left|H_{2}\right|^{2}$ in Eq. (7) may spoil the hierarchy $v_{\phi}\gg v_{EW}$ if their coefficients are of $\mathcal{O}(1)$. However, it is argued in Refs. [65] and [66] that it is natural to set $\lambda_{\phi 1}$ and $\lambda_{\phi 2}$ to be small, since if they vanish, we obtain an enhanced Poincare symmetry for the model. Therefore, the Higgs sector remains radiatively stable even with the spontaneously broken $U(1)_{L}$ at a much higher scale [^1] Without $V_{\mathrm{mix}}$, the scalar sector has a larger symmetry which corresponds to independent rotations of $H_{1}$, $H_{2}$ and $\phi$. If the theory is close to a conformal fixed point below the lepton number breaking scale, the two Higgs doublets are almost massless under this larger symmetry group. The hierarchy $v_{\phi}\gg v_{EW}$ can be equivalently maintained in this way, and the electroweak scale is set by $V_{\mathrm{mix}}$ where $\mu_{12}\sim\mathrm{TeV}$ can be achieved with suitable choices of $\Lambda$, $v_{\phi}$ and $n$..
@@ III Majoron Dark Matter at eV Scales
The Majoron dark matter works much the same as light axionic dark matter, forming a (non-thermal) condensate with the behavior of cold dark matter. We note that when Majoron comprise the dark matter, the previously-discussed natural choice of $F_{J}\gtrsim 10^{11}$GeV corresponds nicely to Majorons in the eV mass range. Beyond the aforementioned theoretical motivations, the eV mass range is important also for practical reasons: 1) using existing telescope observations, we can probe a significant fraction of the parameter space for the Majoron-to-photon decay signals and constrain our Majoron models with $\mathcal{X}_{12}>2$, and 2) for Majorons that can decay into Lyman-Werner photons, the supplied radiation background can be strong enough to trigger the formation of direct collapse black holes [33]. Therefore, this eV mass window could connect the problem of early SMBH formation to the nature of neutrino mass and dark matter, while laying within reach of current instruments like the Hubble Space Telescope [32].
@@ IV SMBH from Majoron Decay
One of the most striking discovery from the JWST is the surprising abundance of active galactic nuclei (AGNs) powered by SMBHs at high redshifts [44, 45, 46, 47, 48, 49, 50, 51, 52]. The existence of this population of seemingly overmassive black holes challenges our understanding of their origin and growth history. Producing such early SMBHs from ‘light seeds’ (e.g. the remnants of Pop III stars) requires extended periods of super-Eddington accretion [83, 84, 85], but even assuming this is possible, feedback events could stall subsequent accretion [86, 87] and it remains challenging to rapidly grow SMBHs from this route. On the other hand, direct collapse black holes [35, 36, 37, 38, 39, 40, 41, 42, 43] can provide heavy seeds for SMBH growth, being born as black holes at masses closer to $10^{5}M_{\odot}$. Furthermore, as estimated in Ref. [88], the comoving number density of the eligible halos for direct collapse has the potential to explain the observed high redshift quasar densities. Although either seed scenarios cannot be excluded by the current JWST data [89, 90], SMBHs discovered in some extreme environments [91, 92, 93] may favor the direct collapse scenario.
@@
$1.\displaystyle k_{\mathrm{H}_{2}}(z)\approx 39\times 0^{-12}\mathrm{~s}^{-1}\frac{J_{LW}(z)}{J_{21}}$.
@@
Due to the in-situ production of radiation, the self-shielding effect should be greatly suppressed and we expect $\varepsilon_{sh}\ll 1$. For a more generic decaying dark matter model, the effect of varying $\varepsilon_{sh}$ was explicitly computed in Ref. [33]. For our more fully realized model here, we use the same chemical network as in Ref. [33] to track the change in temperature and chemical components during the collapse. A successful direct collapse process is identified by a final temperature around $10^{4}~\text{K}$ and an $\mathrm{H_{2}}$ fraction $\ll 10^{-5}$.